# Enhancing the Efficiency of Complex Systems Crystal Structure Prediction by Active Learning Guided Machine Learning Potential

### v1

@@ II.2 Training and refining
When applied to CSP, the training set typically needs to be continuously expanded through the active learning process discussed below. Therefore, ACNN supports restarting training or refining the model from any existing checkpoint with a new or updated dataset. In addition, small perturbations can be introduced to the trainable parameters during refinement to reduce the risk of systematic errors in extrapolated predictions. The underlying intention is for the perturbed model to quickly stabilize in well-sampled regions at the start of refinement, while preserving controlled stochasticity in out-of-distribution regions.
II.3 Active learning
As a data-driven paradigm for interatomic potentials, the performance of ACNN highly depends on the quality of the sampling data. Active learning aims to maintain the predictive accuracy and generalization of the ACNN model during the CSP process, while reducing the number of expensive DFT calculations. According to the principle of maximum information gain, two types of structures are particularly necessary to trigger DFT calculations and should be included in the training set. The first type consists of energy-minimum structures. These samples help reduce the discrepancy between ACNN predictions and DFT reference values, guiding the model toward a more accurate surrogate of the PES. The second stems from unphysical local minima, which typically arise from the model’s insufficient representation of certain regions and therefore require further reinforcement.
The flowchart of the active learning pipeline is illustrated in FIG. 2(a), where two types of data acquisition discussed above are naturally integrated. Upon triggering the workflow, a small set of randomly generated structures undergoes DFT self-consistent field (SCF) calculations, which then form the initial training dataset for ACNN. After training, CSP approach is then carried out on the coarse potential until the number of unphysical structures exceeds a threshold . This indicating that the model’s generalization capability is no longer sufficient for further reliable exploration. The resulting structures are firstly filtered through fingerprint-based[27, 23] deduplication. Subsequently, a thermodynamic phase diagram is constructed based on ACNN accuracy. Despite being incomplete or suboptimal, early ACNN-derived phase diagrams still offer useful guidance for sampling and model refinement. Here, we adopt a composition-wise selection strategy: we first identify the compositions with the lowest convex hull enthalpies (), and then select the lowest-enthalpy structures from each of these compositions. These structures, along with those that fail to converge during optimization, are subsequently subjected to DFT calculations.
The active learning workflow proceeds iteratively, gradually improving the accuracy and generalization of the ACNN model. As its performance advances, more structures are discovered and incorporated into the training set, further reinforcing the model’s capability, as schematically illustrated in FIG. 2(b). Since the sampled dataset also serves as the target for structure prediction, the resulting phase diagram or energy sequence at DFT-level accuracy can be directly constructed from the training data. After sufficient iterations, the phase diagram tends to converge, indicating that the search is approaching exhaustion. Once no additional high value structures are identified, the workflow can be terminated.
Since prediction errors are intrinsic to MLPs, evaluating their influence on the reliability of structure prediction is essential. From the perspective of uncertainty quantification, the error can be decomposed into epistemic and aleatoric components[64]. By systematically sampling underexplored regions of PES, the active learning scheme provides an effective mechanism for reducing epistemic uncertainty, the associated errors can be quantitatively assessed on test sets composed of labeled structures from subsequent iterations. Sources of aleatoric uncertainty mainly contributed by the intrinsic training error of the ACNN approach. Efforts to reduce this uncertainty involve preferentially sampling high-value regions, such as local minima, while avoiding configurations that are far from force equilibrium. Collectively, these errors ultimately manifest as inaccuracies in the predicted energies, forces, and stresses during structure optimization. The energy error, , propagates into the calculation of formation enthalpies, thereby affecting the relative stability of structures within the phase diagram. The force error, , determines the equilibrium configurations obtained via structural optimization and influences the extent of structural distortion. The stress error, , manifests in deviations of lattice parameters and cell geometry, and further impacts the enthalpy through its contribution to the term. If these errors are viewed as perturbations to the equilibrium structures, the resulting phase diagrams tend to overestimate the enthalpy of the structures. To achieve more accurate phase boundaries, it is often necessary to perform full DFT structural relaxations for the structures near the convex hull.
II.4 Computation methods
To assess the effectiveness of the ACNN-based active learning approach in complex material system design, we have conducted systematic tests on two variable composition chemical systems: (i) ternary Mg-Ca-H system, and (ii) quaternary Be-P-N-O system. All known elementary, binary, ternary, and quaternary structures are used as reference structures to evaluate of the thermodynamic stability of newly discovered structures, however, none of them are included a priori in the training set. This ensures a fairer evaluation by more closely reflecting practical application scenarios.
@@ II.4 Computation methods
The ACNN hyperparameter configurations for both systems share the following settings: elemental embedding networks with [64, 64] configuration (two layers of 64 neurons), fitting networks with [256, 256] configuration (two layers of 256 neurons), radial basis functions of order , angular basis functions of order , and an angular momentum cutoff of for characterizing single-bond interactions. The cutoff distances differ between systems: for the Mg-Ca-H system under high pressure, Å and Å are used, whereas for the Be-P-N-O system, Å and Å . In each iteration of the active learning workflow, the ACNN is trained for backpropagation steps. An exponential decay learning rate schedule is adopted, with the learning rate decaying by a factor of 0.97 every steps.
To achieve optimal computational efficiency of ACNN, we employ an in-house implementation of the Broyden–Fletcher–Goldfarb–Shanno (BFGS) algorithm to perform batch structure optimization. The batch mode enables multiple structures to be optimized simultaneously, significantly increasing the degree of parallelism, ensuring efficient utilization of computational resources during the CSP workflow. The convergence criteria were set at an energy threshold of eV/atom and a force threshold of eV/Å for each directional component.
For structure generation, we employ the Crystal Structure Analysis by Particle Swarm Optimization (CALYPSO) code[26, 27] in a purely random sampling mode with only symmetry constraints imposed. This ensures an unbiased evaluation of our approach’s performance by eliminating methodology-specific biases inherent to conventional CSP techniques. Though, the swarm-intelligence-based evolutionary approach has been demonstrated to be highly efficient in practical applications.
@@
III Results and discussion
III.1 Mg-Ca-H system
Recent breakthroughs in superconductivity research have been achieved through the theoretical prediction of clathrate-type hydride compounds under high pressure, such as [37, 38, 39] and [68, 34, 36, 35, 69], which were later confirmed by experimental synthesis. These materials, known as “metal superhydrides”, feature a distinctive structural motif where metal atoms are encapsulated within hydrogen-based clathrate cages. As one of the pioneering examples, exhibits a high superconducting critical temperature () of 215 K at 172 GPa. Theoretical frameworks suggest that lighter metal atoms exhibit lower frequency vibrational modes that favor superconductivity, along with higher electron-phonon coupling from weaker multicentered hydrogen bonding interactions[70]. These insights have motivated the incorporation of Mg into the Ca–H binary system, forming the ternary Mg–Ca–H system as a potential platform for high-temperature superconductivity.
Herein, we conduct a variable stoichiometry structure search at 300 GPa. For the composition space, we restrict the number of Ca and Mg atoms to fewer than 4 each, with the maximum hydrogen-to-metal atomic ratio capped at 16. This resulting in a total of 827 irreducible compositions. Additionally, a constraint of a maximum of 80 atoms per unit cell was imposed, primarily to ensure the inclusion of complex cage-type hydride structures. These settings ensure a sufficiently broad search space, guaranteeing the target structure falls within the search range. Finally, leveraging the high computational efficiency of the ACNN, a total of 5,964,338 structures were optimized during the CSP stage within approximately 2.7 days on a 192-core AMD EPYC 9654 CPU node, achieving a throughput of 474 structures per core-hour.
Following the active learning protocol, DFT SCF calculations were performed on an initial set of 500 randomly generated structures to initiate the computational workflow. The workflow was iteratively executed for a total of 19 cycles. In each iteration, 300 compositions exhibiting the lowest predicted values () were first selected. From each of these compositions, the 3 lowest energy structures () were then chosen as PES sampling candidates for subsequently DFT calculations. In other words, this regulates the computationally expensive yet necessary DFT calculation to a maximum of 900 optimal structures per iteration. The prediction errors for energy, force, and stress, evaluated by comparing the dataset from the -th iteration with the ACNN model trained in the -th iteration, are shown in FIG. 3. This validation scheme quantitatively reflects the performance of the ACNN model in predicting potentially unseen local minima structures. The overall decreasing trend in validation error indicates the model’s improving generalization capability with each iteration. By the final iteration, the ACNN achieved a root mean square error (RMSE) of 44 meV/atom for energy, 283 meV/Å for force, and 14 GPa for cell stresses predictions. This continuous improvement is further evidenced by the progressive increase in the number of uninterrupted structure optimization sequences (as shown in FIG 3(IV)).
Finally, 17, 401 structures were automatically sampled for DFT calculations, including 1, 115 structures necessary to correct unphysical descriptions of the ACNN model, and the remaining 16, 286 nearly local minima structures identified as high value candidates. Among these, 8, 994 structures exhibited energy above the convex hull () below 100 meV/atom. For convex hull refinement, we performed DFT structural relaxation on the 1,000 most stable configurations, with a maximum of approximately 49 meV/atom. As a result, we obtain the phase diagram shown in FIG 4 (a), where a thermodynamically stable Mg2Ca6H46 is successfully located. This structure features a hydrogen network adopting the ‘MEP’ zeolite-type topology [73], with Mg and Ca occupying the 2a and 6d Wyckoff sites, respectively. Alternatively, it can also be viewed as a Mg-substituted variant of the recently reported Ca8H46 phase[74], suggesting its a potential high-temperature superconductor. Here, we found = 137 K at 300 GPa.
Due to the characteristic features of hydride structures, ideal metal superhydride crystal phases are often accompanied by numerous hydrogen-deficient structures, which typically exhibit comparable formation enthalpies. To efficiently expand the pool of candidate clathrate hydrides, we adopt the screening strategy developed by Jiang et al. [75]. This strategy employs an automated topological analysis to identify and classify prototype hydrogen cage motifs. Interestingly, the result accommodates multiple metastable structural prototypes previously reported in zeolite database[76, 77] or other metal hydride systems[78, 79] (see FIG 4 (b)(c)), marked by magenta hollow diamonds in FIG4(a). Nevertheless, we report a new metastable Mg3Ca4H60 phase with an of 66 meV/atom at 300 GPa. In this structure, Mg and Ca atoms occupy the 6b and 8c Wyckoff sites respectively, forming two unusual Mg-centered cage with point group symmetry and Ca-centered cage with symmetry. The H–H bond lengths within these cages measure 0.94, 1.10, 1.11, 1.17 and 1.19 Å at 300 GPa, comparable to those observed in known superconducting hydrogen-rich clathrates, such as CaH6 (1.24 Å at 150GPa)[37] and LaH10 (1.07 and 1.16 Å at 300 GPa)[68, 34].
Building upon the identified candidates structures, we further evaluated their using the SISSO-based prediction model developed by Jiang et al. [75]. This model leveraging the projected density of states of hydrogen at the Fermi level () as an electronic descriptor, in combination with geometric features of the structure facilitates rapid estimation. Benchmarking indicates a mean absolute error of approximately 38 K. The estimated values for these metastable hydrides are presented in FIG5, highlighting the presence of numerous potential high-temperature superconductors that may be experimentally accessible. These results underscore the effectiveness of our algorithm in identifying complex superconducting hydrides with promising properties.
III.2 Be-P-N-O system
In the second test, we apply our method to a more challenging crystal structure prediction (CSP) task by focusing on quaternary compounds. Quaternary system typically present a significantly more complex structural search space due to their greater compositional and configurational degrees of freedom. A particularly promising application lies in the discovery of nonlinear optical (NLO) crystals suitable for deep ultraviolet (DUV) regions. These crystals generate deep ultraviolet coherent light through frequency conversion, enabling essential applications in ultra high resolution lithography, photochemical synthesis, and high precision micromachining. To date, KBe2BO3F2 (KBBF)[15, 80] remains the only practical DUV NLO crystal. However, its strong layering habit[81] has significantly hindered its further commercialization. In contrast, phosphates exhibit diverse non-centrosymmetric (NCS) structures, short UV cutoff edges, and favorable single-crystal growth behavior, making them as strong candidates for the discovery of new NLO materials[82, 83, 84]. Given these advantages, the Be-P-N-O system is deemed a suitable test case for evaluating our approach.
A total of 9,381,987 structures were fully relaxed during the CSP stage by ACNN model, executed over 29 active learning iterations to efficiently explore the complex configurational landscape. The entire relaxation process took approximately 1.8 days on a 192-core AMD 9654 CPU, achieving a speed of about 1131 structures per core-hour. Improved relaxation efficiency relative to the previous ternary system primarily attributed to the reduced average number of atoms per structure, as well as the ambient pressure conditions and larger atomic radii (relative to hydrogen) decrease the number of neighboring atoms, collectively lowering the computational workload. For structure sampling, a more flexible strategy was employed. During the initial five iterations, two distinct structures with the lowest energy () were selected from each of the 270 compositions () with the lowest predicted , aiming to facilitate broader exploration of the compositional space. In subsequent iterations, the sampling parameters were adjusted to and to more densely probe novel structures within the compositional region that had been largely validated as near the convex hull and potentially thermodynamical stable.
The corresponding prediction errors during each iterations are presented in FIG6(a). The quaternary system exhibits a more gradual decline in validation error, consistent with the observed trend reported by Wang et al. [42] when transitioning from binary to ternary systems. This slower convergence likely arises from the increased structural complexity of higher-order systems, which imposes greater demands on both CSP and the generalization capacity of the MLP. The ACNN model ultimately achieves a RMSE of 62 meV/atom for energy, 325 meV/Å for forces, and 2.2 GPa for cell stresses predictions.
A total of 13,214 structures were selected for SCF calculations, comprising 11,460 local minima and 1,754 unphysical configurations mistakenly predicted as low-enthalpy by ACNN. Among these, top 1,000 structures with the lowest values were subjected to full DFT structural optimization to construct a more accurate convex hull surface. The complete quaternary phase diagram is presented in the Supplementary Information. Three thermodynamically stable phases were identified: -BePNO2, -BeP3NO7, and -BeP3N5O. These phases are distributed within the pseudo-ternary compositional space spanned by BeO-P2O5-P3N5, as shown in FIG 6(b). As illustrated in FIG 6(c), the crystal structures of all three phases are composed of P-centered and Be-centered tetrahedra, interconnected through corner sharing to form robust three-dimensional (3D) frameworks with distinctive cavity geometries. This structure structural motif resembles those reported in prior studies[85, 86], where nitrogen atoms partially substitute oxygen within phosphate frameworks, forming PNxO4-x (x = 1, 2, 3, 4) mixed anionic tetrahedra. Such hybrid coordination environments contribute to the structural diversity and potential functionality of these materials.
The electronic band structures of the thermodynamically stable phases are presented in FIG.7, revealing band gaps of 5.85 eV (212 nm), 5.56 (223 nm), and 3.3 (376 nm), as calculated at the PBE level. While it is well-established that PBE tends to underestimate band gaps, the first two phases remain within the DUV spectral range, suggesting their potential for DUV optical applications. Additionally, the resulting structure pool include nearly 3,000 structures with formation enthalpies within , commonly associated with potential experimental synthesizability. A more comprehensive analysis of the physical properties of these metastable candidates will provided in a separate forthcoming work.
IV Conclusions
In summary, we present ACNN integrated with the CSP technique to establish an efficient and robust active learning workflow aimed at addressing the high computational complexity inherent in predicting complex multicomponent systems. The robust active learning scheme and the effectiveness of the ACNN leverages the ability to explore millions of candidate structures across various stoichiometries and large atomic configurations, enable a more comprehensive investigation of complex systems. We first apply the algorithm to the high pressure Mg–Ca–H system as a test case, where we discover a stable phase along with a series of metastable candidates as promising high- superconductors. We further demonstrate the effectiveness of our method on the more complex quaternary Be–P–N–O system. In this case, we identify two potential DUV NLO crystals: -BePNO2 and -BeP3NO7. Both cases not only corroborate the rich diversity of promising compounds in complex multicomponent systems, but also highlight the capability of our algorithm to efficiently discover functional materials within vast and high-dimensional chemical spaces.
The modular design of the active learning framework offers significant flexibility for application and external integration. This enables dynamic adjustments to the sampling strategy and search space during runtime, as demonstrated in the quaternary Be–P–N–O case above. The ACNN software package is released alongside the first-principles simulation code ARES, forming a technically integrated and robust software ecosystem. By facilitating software–hardware co-design and optimization, this integration fully unleashes the potential of ACNN for materials design and downstream property simulations. Such software design principles provide a computational foundation for automated materials discovery.
@@ Acknowledgements.
This work is supported by the National Key R&D Program of China (Grant No. 2023YFB3003001), the National Natural Science Foundation of China (Grant No. 12374008 and 12022408), the Interdisciplinary Integration and Innovation Project of JLU, and Fundamental Research Funds for the Central Universities and the Program for JLU Science and Technology Innovative Research Team (JLUSTIRT). We also thank the Changchun Computing Center and the Eco-Innovation Center for providing comprehensive computing resources and technical support throughout the completion of this work.References
- Green et al. [2014] M. A. Green, A. Ho-Baillie, and H. J. Snaith, The emergence of perovskite solar cells, Nature photonics 8, 506 (2014).
@@
- Sun et al. [2023] H. Sun, M. Huo, X. Hu, J. Li, Z. Liu, Y. Han, L. Tang, Z. Mao, P. Yang, B. Wang, et al., Signatures of superconductivity near 80 k in a nickelate under high pressure, Nature 621, 493 (2023).
@@
- Hart et al. [2021] G. L. Hart, T. Mueller, C. Toher, and S. Curtarolo, Machine learning for alloys, Nature Reviews Materials 6, 730 (2021).
- Tsai and Yeh [2014] M.-H. Tsai and J.-W. Yeh, High-entropy alloys: a critical review, Materials Research Letters 2, 107 (2014).
@@
- Chen et al. [1989] C. Chen, Y. Wu, A. Jiang, B. Wu, G. You, R. Li, and S. Lin, New nonlinear-optical crystal: Lib3o5, Journal of the Optical Society of America B 6, 616 (1989).
- Nikogosyan [1991] D. Nikogosyan, Beta barium borate (bbo) a review of its properties and applications, Applied Physics A 52, 359 (1991).
@@
- Oganov and Valle [2009] A. R. Oganov and M. Valle, How to quantify energy landscapes of solids, The Journal of chemical physics 130 (2009).
- Villars and Iwata [2013] P. Villars and S. Iwata, Pauling file verifies/reveals 12 principles in materials science supporting four cornerstones given by nature, Chemistry of metals and alloys , 81 (2013).
- Butler et al. [2016] K. T. Butler, J. M. Frost, J. M. Skelton, K. L. Svane, and A. Walsh, Computational materials design of crystalline solids, Chemical Society Reviews 45, 6138 (2016).
- Pickard and Needs [2006] C. J. Pickard and R. Needs, High-pressure phases of silane, Physical review letters 97, 045504 (2006).
@@
- Oganov et al. [2019] A. R. Oganov, C. J. Pickard, Q. Zhu, and R. J. Needs, Structure prediction drives materials discovery, Nature Reviews Materials 4, 331 (2019).
@@
- Shi et al. [2018a] X. Shi, C. He, C. J. Pickard, C. Tang, and J. Zhong, Stochastic generation of complex crystal structures combining group and graph theory with application to carbon, Phys. Rev. B 97, 014104 (2018a).
- Wang et al. [2023] J. Wang, H. Gao, Y. Han, C. Ding, S. Pan, Y. Wang, Q. Jia, H.-T. Wang, D. Xing, and J. Sun, Magus: machine learning and graph theory assisted universal structure searcher, National Science Review 10, nwad128 (2023).
- Zhu et al. [2022] Q. Zhu, B. Kang, and K. Parrish, Symmetry relation database and its application to ferroelectric materials discovery, MRS Communications 12, 686 (2022).
- Fredericks et al. [2021] S. Fredericks, K. Parrish, D. Sayre, and Q. Zhu, Pyxtal: A python library for crystal structure generation and symmetry analysis, Computer Physics Communications 261, 107810 (2021).
- Shi et al. [2018b] X. Shi, C. He, C. J. Pickard, C. Tang, and J. Zhong, Stochastic generation of complex crystal structures combining group and graph theory with application to carbon, Physical Review B 97, 014104 (2018b).
- Ma et al. [2009] Y. Ma, M. Eremets, A. R. Oganov, Y. Xie, I. Trojan, S. Medvedev, A. O. Lyakhov, M. Valle, and V. Prakapenka, Transparent dense sodium, Nature 458, 182 (2009).
@@
- Somayazulu et al. [2019] M. Somayazulu, M. Ahart, A. K. Mishra, Z. M. Geballe, M. Baldini, Y. Meng, V. V. Struzhkin, and R. J. Hemley, Evidence for superconductivity above 260 k in lanthanum superhydride at megabar pressures, Physical review letters 122, 027001 (2019).
- Li et al. [2022] Z. Li, X. He, C. Zhang, X. Wang, S. Zhang, Y. Jia, S. Feng, K. Lu, J. Zhao, J. Zhang, et al., Superconductivity above 200 k discovered in superhydrides of calcium, Nature communications 13, 2863 (2022).
- Deringer et al. [2018] V. L. Deringer, C. J. Pickard, and G. Csányi, Data-driven learning of total and local energies in elemental boron, Physical review letters 120, 156001 (2018).
- Podryabinkin et al. [2019] E. V. Podryabinkin, E. V. Tikhonov, A. V. Shapeev, and A. R. Oganov, Accelerating crystal structure prediction by machine-learning interatomic potentials with active learning, Physical Review B 99, 064114 (2019).
- Merchant et al. [2023] A. Merchant, S. Batzner, S. S. Schoenholz, M. Aykol, G. Cheon, and E. D. Cubuk, Scaling deep learning for materials discovery, Nature 624, 80 (2023).
- Batatia et al. [2023] I. Batatia, P. Benner, Y. Chiang, A. M. Elena, D. P. Kovács, J. Riebesell, X. R. Advincula, M. Asta, W. J. Baldwin, N. Bernstein, et al., A foundation model for atomistic materials chemistry, arXiv preprint arXiv:2401.00096 (2023).
- Yang et al. [2024] H. Yang, C. Hu, Y. Zhou, X. Liu, Y. Shi, J. Li, G. Li, Z. Chen, S. Chen, C. Zeni, et al., Mattersim: A deep learning atomistic model across elements, temperatures and pressures, arXiv preprint arXiv:2405.04967 (2024).
- Xie et al. [2024] F. Xie, T. Lu, S. Meng, and M. Liu, Gptff: A high-accuracy out-of-the-box universal ai force field for arbitrary inorganic materials, Science Bulletin 69, 3525 (2024).
- Deng et al. [2023] B. Deng, P. Zhong, K. Jun, J. Riebesell, K. Han, C. J. Bartel, and G. Ceder, Chgnet as a pretrained universal neural network potential for charge-informed atomistic modelling, Nature Machine Intelligence 5, 1031 (2023).
- Zhang et al. [2024] D. Zhang, X. Liu, X. Zhang, C. Zhang, C. Cai, H. Bi, Y. Du, X. Qin, A. Peng, J. Huang, et al., Dpa-2: a large atomic model as a multi-task learner, npj Computational Materials 10, 293 (2024).
- Wang et al. [2025] X. Wang, C. Zhang, Z. Wang, H. Liu, J. Lv, H. Wang, Y. Ma, et al., Discovery of high-temperature superconducting ternary hydrides via deep learning, arXiv preprint arXiv:2502.16558 (2025).
- Behler [2011] J. Behler, Atom-centered symmetry functions for constructing high-dimensional neural network potentials, The Journal of chemical physics 134 (2011).
- Bartók et al. [2013] A. P. Bartók, R. Kondor, and G. Csányi, On representing chemical environments, Physical Review B—Condensed Matter and Materials Physics 87, 184115 (2013).
- Hong et al. [2020] F. Hong, L. Yang, P. Shan, P. Yang, Z. Liu, J. Sun, Y. Yin, X. Yu, J. Cheng, and Z. Zhao, Superconductivity of lanthanum superhydride investigated using the standard four-probe configuration under high pressures, Chinese physics letters 37, 107401 (2020).
@@
- Sun et al. [2019] Y. Sun, J. Lv, Y. Xie, H. Liu, and Y. Ma, Route to a superconducting phase above room temperature in electron-doped hydride compounds under high pressure, Physical review letters 123, 097001 (2019).
@@
- Chen et al. [2009] C. Chen, G. Wang, X. Wang, and Z. Xu, Deep-uv nonlinear optical crystal kbe 2 bo 3 f 2—discovery, growth, optical properties and applications, Applied Physics B 97, 9 (2009).
- Peng et al. [2018] G. Peng, N. Ye, Z. Lin, L. Kang, S. Pan, M. Zhang, C. Lin, X. Long, M. Luo, Y. Chen, et al., Nh4be2bo3f2 and -be2bo3f: Overcoming the layering habit in kbe2bo3f2 for the next-generation deep-ultraviolet nonlinear optical materials, Angewandte Chemie 130, 9106 (2018).
- Zhao et al. [2014] S. Zhao, P. Gong, S. Luo, L. Bai, Z. Lin, C. Ji, T. Chen, M. Hong, and J. Luo, Deep-ultraviolet transparent phosphates rbba2 (po3) 5 and rb2ba3 (p2o7) 2 show nonlinear optical activity from condensation of [po4] 3–units, Journal of the American Chemical Society 136, 8560 (2014).
@@
- Jin et al. [2023] W. Jin, C. Xie, X. Hou, M. Cheng, E. Tikhonov, M. Wu, S. Pan, and Z. Yang, From monofluorophosphates a2po3f to difluorophosphates apo2f2 (a= alkali metal): Design of a potential deep-ultraviolet nonlinear optical materials system with a shortened phase-matching wavelength, Chemistry of Materials 35, 5281 (2023).

### v2

@@ IV.1 Training and refining protocol
When applied to CSP, the training set typically needs to be continuously expanded through the self-optimizing process discussed below. ACNN supports restarting training or refining the model from any existing checkpoint with a new or updated dataset. During refinement, small perturbations are introduced to the trainable parameters, which helps mitigate systematic errors in extrapolated predictions. The rationale is that the perturbed model can quickly stabilize in well-sampled regions at the onset of refinement, while maintaining controlled stochasticity in out-of-distribution regions.
@@
The ACNN hyperparameter configurations for both systems share the following settings: elemental embedding networks with [64, 64] configuration (two layers of 64 neurons), fitting networks with [256, 256] configuration (two layers of 256 neurons), radial basis functions of order , angular basis functions of order , and an angular momentum cutoff of for characterizing single-bond interactions. The cutoff distances differ between systems: for the Mg-Ca-H system under high pressure, Å and Å are used, whereas for the Be-P-N-O system, Å and Å . In each iteration of the self-optimizing workflow, the ACNN is trained for backpropagation steps. During training, an exponential decay learning rate schedule is adopted, where the learning rate decays by a factor of 0.97 every steps. The loss prefactors , , and were initially set to 0.1, 10 and 0.1, respectively. While and were kept constant, was exponentially decayed to 0.1 using the same decay factor of 0.97 every steps.
IV.2 Structure generation
Structure generation was performed using the Crystal Structure Analysis by Particle Swarm Optimization (CALYPSO) code[17, 18] in a purely random sampling mode, applying only stochastic symmetry constraints. This approach enables an unbiased assessment of our method by deliberately excluding the global structural evolution operators commonly integrated into conventional CSP workflows, which, although often statistically efficient, can introduce method-dependent biases.
Minimum interatomic distances and cell volumes are specified to prevent the generation of unphysical configurations at the outset. For the Mg–Ca–H system, the initial minimum distances (Å) are given by the symmetric matrix,
where the matrix is symmetric with rows and columns ordered as (Mg, Ca, H). The initial cell volume is estimated as a linear combination of atomic contributions,
For the Be–P–N–O system, the initial minimum distances (Å) are given by,
with rows and columns ordered as (Be, P, N, O). The atomic volumes are
IV.3 Calculation details
@@
To achieve optimal computational efficiency of ACNN, we employ an in-house implementation of the Broyden–Fletcher–Goldfarb–Shanno (BFGS) algorithm [67] to perform batch structure optimization. Atomic positions and lattice parameters are optimized using Hellmann–Feynman forces and the computed stress tensor, respectively, both derived from the ACNN potential. Convergence was achieved when the enthalpy change fell below eV/atom and the maximum atomic force was less than eV/Å . Here, the batch mode enables multiple structures to be optimized simultaneously, significantly increasing the degree of parallelism.
Competing Interests
The authors declare no conflict of interest.
Data Availability
Some data supporting the findings of this study are available in the Supplementary Materials. Additional data are available from the corresponding author upon reasonable request.
Code Availability
The custom code developed and used in this study will be made publicly available upon publication. Prior to that, it can be obtained from the corresponding author upon reasonable request.
Author contributions
Y.X., H.L., C.X., and J.L. designed the research. J.L., J.L., and X.Z. developed the code. J.L. conducted the majority of the CSP calculations. J.L., Q.S., J.F., and B.J. performed the property calculations. J.L., J.L., H.L., K.B., C.X., Y.X., and Y.M. analyzed and interpreted the data and contributed to writing the manuscript.
@@ Acknowledgements.
This work is supported by the National Key R&D Program of China (Grant No. 2023YFB3003001), the National Natural Science Foundation of China (Grant No. 12374008 and 12022408), the National Natural Science Foundation of China (52403305), the Strategic Priority Research Program of the Chinese Academy of Sciences (XDB0880000), the Interdisciplinary Integration and Innovation Project of JLU, the Fundamental Research Funds for the Central Universities and the Program for JLU Science and Technology Innovative Research Team (JLUSTIRT). We also thank the Changchun Computing Center and the Eco-Innovation Center for providing comprehensive computing resources and technical support throughout the completion of this work.References
@@
- Otto et al. [2013] F. Otto, A. Dlouhỳ, C. Somsen, H. Bei, G. Eggeler, and E. P. George, The influences of temperature and microstructure on the tensile properties of a cocrfemnni high-entropy alloy, Acta Materialia 61, 5743 (2013).
@@
- Wu et al. [1996] B. Wu, D. Tang, N. Ye, and C. Chen, Linear and nonlinear optical properties of the kbe2bo3f2 (kbbf) crystal, Optical Materials 5, 105 (1996).
@@
- Park et al. [2025] H. Park, A. Onwuli, K. T. Butler, and A. Walsh, Mapping inorganic crystal chemical space, Faraday Discussions 256, 601 (2025).
@@
- Batatia et al. [2022] I. Batatia, D. P. Kovacs, G. Simm, C. Ortner, and G. Csányi, Mace: Higher order equivariant message passing neural networks for fast and accurate force fields, Advances in Neural Information Processing Systems 35, 11423 (2022).
- Juelsholt [2025] M. Juelsholt, Continued challenges in high-throughput materials predictions: Mattergen predicts compounds from the training dataset., chemrxiv-2025-mkls8 (2025).
- Tong et al. [2018] Q. Tong, L. Xue, J. Lv, Y. Wang, and Y. Ma, Accelerating calypso structure prediction by data-driven learning of a potential energy surface, Faraday discussions 211, 31 (2018).
- Tong et al. [2020] Q. Tong, P. Gao, H. Liu, Y. Xie, J. Lv, Y. Wang, and J. Zhao, Combining machine learning potential and structure prediction for accelerated materials design and discovery, The journal of physical chemistry letters 11, 8710 (2020).
- Curtarolo et al. [2012] S. Curtarolo, W. Setyawan, G. L. Hart, M. Jahnatek, R. V. Chepulskii, R. H. Taylor, S. Wang, J. Xue, K. Yang, O. Levy, et al., Aflow: An automatic framework for high-throughput materials discovery, Computational Materials Science 58, 218 (2012).
- Pizzi et al. [2016] G. Pizzi, A. Cepellotti, R. Sabatini, N. Marzari, and B. Kozinsky, Aiida: automated interactive infrastructure and database for computational science, Computational Materials Science 111, 218 (2016).
- Yu et al. [2024] H. Yu, M. Giantomassi, G. Materzanini, J. Wang, and G.-M. Rignanese, Systematic assessment of various universal machine-learning interatomic potentials, Materials Genome Engineering Advances 2, e58 (2024).
- Liu et al. [2021] X. Liu, H. Niu, and A. R. Oganov, Copex: co-evolutionary crystal structure prediction algorithm for complex systems, npj Computational Materials 7, 199 (2021).
@@
- Lyakhov et al. [2010] A. O. Lyakhov, A. R. Oganov, and M. Valle, How to predict very large and complex crystal structures, Computer Physics Communications 181, 1623 (2010).
- Antunes et al. [2024] L. M. Antunes, K. T. Butler, and R. Grau-Crespo, Crystal structure generation with autoregressive large language modeling, Nature Communications 15, 1 (2024).
- Luo et al. [2024] X. Luo, Z. Wang, P. Gao, J. Lv, Y. Wang, C. Chen, and Y. Ma, Deep learning generative model for crystal structure prediction, npj Computational Materials 10, 254 (2024).
- Xie et al. [2021] T. Xie, X. Fu, O.-E. Ganea, R. Barzilay, and T. Jaakkola, Crystal diffusion variational autoencoder for periodic material generation, arXiv preprint arXiv:2110.06197 (2021).
- Vasylenko et al. [2024] A. Vasylenko, B. M. Asher, C. M. Collins, M. W. Gaultois, G. R. Darling, M. S. Dyer, and M. J. Rosseinsky, Inferring energy–composition relationships with bayesian optimization enhances exploration of inorganic materials, The Journal of chemical physics 160 (2024).
- Nocedal and Wright [2006] J. Nocedal and S. J. Wright, Numerical Optimization, 2nd ed. (Springer, New York, NY, USA, 2006).